# Accelerating Focal Search in Multi-Agent Path Finding with Tighter Lower Bounds

### v1

@@ V-B Performance
We first show that double search reduces the number of low-level nodes in FOCAL and high-level CT nodes. As shown in Figure 2, each data point in the figure represents a test case that both DECBS and ECBS completed within the time limit. On average, DECBS only searches 212k nodes in the low-level FOCAL compared to ECBS’s 416k nodes, achieving nearly 50% node savings. For high-level CT nodes, DECBS expands only 291 nodes compared to 427 for ECBS, resulting in nearly 30% node savings. Although DECBS performs an additional shortest path search in the low-level search, Figure 3 (columns 2 and 4) shows that the total runtime of DECBS for test cases is lower than that of ECBS, indicating that the time saved by reducing FOCAL nodes and CT nodes more than compensates for the extra shortest path search cost, especially in high agent density scenarios.
@@
Figure 3 shows detailed success rates over different maps, numbers of agents, suboptimality factors, and various optimization methods. In the column 1 and 3 subfigures of Figure 3, we can observe that DECBS outperforms ECBS in success rate on all small and medium maps. In particular, on the warehouse map, there is a noticeable gap between the green line for DECBS+BC+TR and the other lines. Also, on the den_312d map with $w=1.1$, DECBS outperforms ECBS+BC+TR, and all three variants—DECBS, DECBS+BC, and DECBS+BC+TR—consistently outperform the other methods under different conditions. As for the column 2 and 4 subfigures of Figure 3, we plot the runtimes of the two algorithms using different colors to indicate the number of agents. It is clear that as the number of agents increases, the acceleration effect becomes more pronounced. As agent density grows, it becomes harder for focal search to find a low-cost goal state. In contrast, with double search, we obtain the highest LB value, which provides a better chance to jump over unpromising nodes and devote more time to nodes with fewer conflicts.
@@
However, in Figure 3, on large maps such as den_520d and Paris_1_256, we can observe that DECBS is slower than ECBS. This is mainly because, compared to the map sizes, the number of agents is too small. With such low agent density, collisions between agents rarely occur. In this case, focal search behaves more like $A^{*}$ search, and double search essentially becomes two $A^{*}$ searches. However, we can also observe in the runtime subfigures that as the number of agents increases, the crossover point moves further and further above the 1x line.
@@
We present the average speedup of DECBS over ECBS for different agent numbers and suboptimality factors in Figure 4. Overall, DECBS is 23.5% faster than ECBS on average with BC+TR optimization in small and medium maps, and as the number of agents increases, the improvements become more pronounced. Moreover, when both algorithms use the same optimizations (such as BC+TR), the optimization method yields even more performance gains for DECBS. Also, as the suboptimality factor increases, DECBS exhibits more improvements compared to ECBS. However, on large maps, DECBS is slower than ECBS. Still, we observe that as the number of agents increases, the gap between DECBS and ECBS shrinks. These large map results support our earlier statements: in large maps with low agent density, focal search degenerates into $A^{*}$ search, which means DECBS essentially performs two $A^{*}$ searches. That is why, when the number of agents increases, the speed of DECBS improves and approaches that of ECBS because collisions become more frequent. Regarding suboptimality, the performance degradation remains similar because, in large maps, there are hardly any collisions between agents. In such cases, the shortest path is likely collision-free, making the FOCAL search heuristic ineffective.

### v2

@@ V-B Performance
We first show double search reduces the number of low-level nodes in FOCAL and high-level CT nodes. As shown in Figure˜2, each data point in the figure represents a test case that both DECBS and ECBS completed within the time limit. On average, DECBS only searches 212k nodes in the low-level FOCAL compared to ECBS’s 416k nodes, achieving nearly 50% node savings. For high-level CT nodes, DECBS expands only 291 nodes compared to 427 for ECBS, resulting in nearly 30% node savings. Although DECBS performs an additional shortest path search in the low-level search, Figure˜3 (columns 2 and 4) shows that the total runtime of DECBS for test cases is lower than that of ECBS, indicating that the time saved by reducing FOCAL nodes and CT nodes more than compensates for the extra shortest path search cost, especially in high agent density scenarios.
@@
Figure˜3 shows detailed success rates over different maps, numbers of agents, suboptimality factors, and various optimization methods. In the column 1 and 3 subfigures of Figure˜3, we can observe that DECBS outperforms ECBS in success rate on all small and medium maps. In particular, on the warehouse map, there is a noticeable gap between the green line for DECBS+BC+TR and the other lines. Also, on the den_312d map with $w=1.1$, DECBS outperforms ECBS+BC+TR, and all three variants—DECBS, DECBS+BC, and DECBS+BC+TR—consistently outperform the other methods under different conditions. As for the column 2 and 4 subfigures of Figure˜3, we plot the runtimes of the two algorithms using different colors to indicate the number of agents. It is clear that as the number of agents increases, the acceleration effect becomes more pronounced. As agent density grows, it becomes harder for focal search to find a low-cost goal state. In contrast, with double search, we obtain the highest LB value, which provides a better chance to jump over unpromising nodes and devote more time to nodes with fewer conflicts.
@@
However, in Figure˜3, on large maps such as den_520d and Paris_1_256, we can observe that DECBS is slower than ECBS. This is mainly because, compared to the map sizes, the number of agents is too small. With such low agent density, collisions between agents rarely occur. In this case, focal search behaves more like $A^{*}$ search, and double search essentially becomes two $A^{*}$ searches. However, we can also observe in the runtime subfigures that as the number of agents increases, the crossover point moves further and further above the 1x line.
@@
We present the average speedup of DECBS over ECBS for different agent numbers and suboptimality factors in Figure˜4. Overall, DECBS is 23.5% faster than ECBS on average with BC+TR optimization in small and medium maps, and as the number of agents increases, the improvements become more pronounced. Moreover, when both algorithms use the same optimizations (such as BC+TR), the optimization method yields even more performance gains for DECBS. Also, as the suboptimality factor increases, DECBS exhibits more improvements compared to ECBS. However, on large maps, DECBS is slower than ECBS. Still, we observe that as the number of agents increases, the gap between DECBS and ECBS shrinks. These large map results support our earlier statements: in large maps with low agent density, focal search degenerates into $A^{*}$ search, which means DECBS essentially performs two $A^{*}$ searches. That is why, when the number of agents increases, the speed of DECBS improves and approaches that of ECBS because collisions become more frequent. Regarding suboptimality, the performance degradation remains similar because, in large maps, there are hardly any collisions between agents. In such cases, the shortest path is likely collision-free, making the FOCAL search heuristic ineffective.
@@ VI Conclusion
VII Acknowledgement
The research at the University of California, Irvine, the Carnegie Mellon University and the University of Southern California was supported by the National Science Foundation (NSF) under grant numbers 2328671, 2441629, 2434916, 2321786, 2112533, and 2121028, as well as gifts from Amazon Robotics and the Donald Bren Foundation. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the sponsoring organizations, agencies, or the U.S. government.